OpenStudy (x3_drummerchick):

help! WILL GIVE MEDALS!! i have no idea how to start this... find sin2x, cos2x, and tan2x given cscx=5, tanx<0

OpenStudy (welshfella):

tan x is positive in Q III

OpenStudy (x3_drummerchick):

imm sorry, i put the wrong problem up

OpenStudy (x3_drummerchick):

its supposed to be CSCx=5, tanx<0

OpenStudy (welshfella):

ok then that angle is in the second quadrant csc x = 1/ sin x sin x = 1/5

OpenStudy (x3_drummerchick):

ok, so then do i just do the formulas again, use cos2x and find sin2x?

OpenStudy (welshfella):



OpenStudy (x3_drummerchick):

cos2a= \[1-2\sin ^{2}x\]

OpenStudy (x3_drummerchick):

ohh wait so do the opposite over hypotenuse thing?

OpenStudy (welshfella):

you can find cos x , tan x from the above digram

OpenStudy (welshfella):

yes